# A Guide to Misinformation Detection Data and Evaluation

### v2

@@ Abstract.
Misinformation is a complex societal issue, and mitigating solutions are difficult to create due to data deficiencies. To address this, we have curated the largest collection of (mis)information datasets in the literature, totaling 75. From these, we evaluated the quality of 36 datasets that consist of statements or claims, as well as the 9 datasets that consist of data in purely paragraph form. We assess these datasets to identify those with solid foundations for empirical work and those with flaws that could result in misleading and non-generalizable results, such as spurious correlations, or examples that are ambiguous or otherwise impossible to assess for veracity. We find the latter issue is particularly severe and affects most datasets in the literature. We further provide state-of-the-art baselines on all these datasets, but show that regardless of label quality, categorical labels may no longer give an accurate evaluation of detection model performance. Finally, we propose and highlight Evaluation Quality Assessment (EQA) as a tool to guide the field toward systemic solutions rather than inadvertently propagating issues in evaluation. Overall, this guide aims to provide a roadmap for higher quality data and better grounded evaluations, ultimately improving research in misinformation detection. All datasets and other artifacts are available at misinfo-datasets.complexdatalab.com.
@@ 1. Introduction
Finally, our findings come together in our recommendation for a new practice of Evaluation Quality Assessment (EQA), which offers a tool to improve data selection and evaluation methodologies. In summary, we present a guide to misinformation datasets, including:
@@
  • Recommended practices of EQA, for research proposing misinformation detection methods. Informed by our findings of severe quality issues in both data and evaluation procedures, such as how simple metrics like accuracy and F1 may be obsolete in this domain, we propose that assessment of evaluation quality is an indispensable component for future research here.
@@ 3. Data Quality
Shortcut learning is a serious barrier to predictive systems working in the real world (Geirhos et al., 2020). In this section, we assess datasets’ potential for teaching algorithms spurious keyword and temporal correlations. We then assess whether the inputs are making sufficiently complete and unambiguous claims for it to be feasible to predict their veracity at all. We provide the automated tools that enable this analysis through our website as the Complex Data Lab Dataset Quality Assessment (CDL-DQA).
@@ 3.3. Feasibility
We first aggregate over datasets, presenting the individual results of our 9 human and AI annotators in Figure 4. Our results show universal agreement that without an evidence retrieval (search) system, at least half of claims in these 29 datasets cannot be validly assessed for veracity. This suggests that such methods are often being evaluated on impossible tasks, and there is a severe risk that evaluation comparing such methods is determining not the best generalizable predictor but the best shortcut learner. We recommend that any proposed methods in this category conduct stringent Evaluation Quality Assessment (see Section 5), and suggest that applications requiring even slightly out of distribution predictions should ground them in evidence.
@@
This has 3 main implications. First, there is an urgent need for higher quality claims datasets. Second, supplementary information beyond claim text, such as claim dates, authors, and other additional context may help alleviate this problem. However, it is critical to assess how the specific information a predictive method processes impacts the feasibility of the data it is being trained and evaluated on, to confirm that performance margins between methods represent real progress and not progress in predicting noise. Thus, third, we propose that even in the more favorable retrieval-augmented setting, an Evaluation Quality Assessment (Section 5) is paramount.
@@ 4.2. The Flaw in Current Metrics
Dataset Label $\neq$ Prediction Label $=$ Prediction
Rationale is Valid Invalid Valid Invalid
@@
5. Evaluation Quality Assessment
@@
We have highlighted multiple ways datasets and evaluation procedures could deliver results that will fail to translate to the real world. While the information here directly reveals some pitfalls, we suspect that future research—with new datasets, new methods, and new needs—will encounter similar problems. To improve the validity of conclusions in our field in a systematic and future-proof way, we propose the Evaluation Quality Assessment (EQA) as a key component of an improved research methodology.
@@
We draw inspiration from the common practices for systematically and transparently documenting dataset publications, such as “datasheets for datasets” (Gebru et al., 2021), and suggest that similar attention should be given to the usage of datasets. However, it is important to ensure data and evaluation quality not just as a lengthy post-hoc appendix, but as an integral part of the research process and a key foundation of every paper’s core arguments. Therefore, we propose including EQA as a scientifically mandatory step within the methodology. This means augmenting the standard descriptive discussion of data used with a critical analysis of it, explaining what steps and quality assessments were conducted to assure that the data is suitable to prove the experimental conclusions. We suggest that dedicating even a paragraph to this, rather than taking quality of “standard” datasets for granted, could dramatically improve the validity of our empirical results and the rate of real progress in the field. Conversely, current practices risk misleading evaluations, generalization failures, and immeasurable time and effort dedicated to measurements of noise.
We make three specific recommendations for Evaluation Quality Assessment as follows.
  • Papers should express clearly the bounds on the intended generalization of their method, and in turn what types of generalization the evaluation aims to prove. For example, at one extreme, if the evaluation data is a perfect IID sample of the application data and there is no distribution shift whatsoever, then many spurious correlations and other quality issues are irrelevant. But most methods are intended to apply to data with distributional differences in time, topic, source, or other characteristics. By briefly but concretely laying out which distributional shifts are in scope and which are not in scope, a paper states their evaluation hypotheses.
@@
  • Papers should report analyses performed on the data and evaluation process to assure it is suitable to test the intended generalization of the method. One universal baseline is that author should qualitatively inspect at minimum 25 random examples—input, output, and ground truth—where their method made correct predictions, and 25 where it made incorrect ones. This data should be provided for reproducibility, and provides a basic sanity check. We note, for example, that the feasibility and categorical metrics issues we have reported would be detected by such a sanity check in the majority of datasets we examined. Furthermore, papers should also perform quantitive assessments of their data and evaluation, such as (but not limited to) the ones discussed in prior sections and available in our CDL-DQA toolkit.
  • Papers should report limitations in EQA either directly in methodology or in the standard limitations section. Authors should consider not just surface-level limitations like the number or size of datasets in the experiments, but also potential evaluation quality issues that the EQA could not rule out. For instance, if generalization under temporal distribution shifts is expected but the method was trained on data not checked for spurious temporal correlations, this would represent a limitation that future research should be aware of and consider addressing.
@@
In this way, we treat assessing evaluation as a crucial step for validating the research process. For authors, this will help reinforce the demonstrated rigour of the experiments and the validity of proposed methods. It will also prevent the rejection of strong methods due to unrealistically high baseline settings caused by spurious correlations or other evaluation issues. For reviewers, this will produce more explicit grounding (or lack thereof) for the experiments, as well as a framework to recommend when evaluation is insufficiently justified and needs better validation. Thus, we propose that adopting EQA can efficiently and substantively mitigate propagating and compounding errors arising in this domain from current practices of copy-pasting evaluation from past literature.

### v3

@@ Abstract.
Misinformation is a complex societal issue, and mitigating solutions are difficult to create due to data deficiencies. To address this, we have curated the largest collection of (mis)information datasets in the literature, totaling 75. From these, we evaluated the quality of 36 datasets that consist of statements or claims, as well as the 9 datasets that consist of data in purely paragraph form. We assess these datasets to identify those with solid foundations for empirical work and those with flaws that could result in misleading and non-generalizable results, such as spurious correlations, or examples that are ambiguous or otherwise impossible to assess for veracity. We find the latter issue is particularly severe and affects most datasets in the literature. We further provide state-of-the-art baselines on all these datasets, but show that regardless of label quality, categorical labels may no longer give an accurate evaluation of detection model performance. Finally, we propose and highlight Evaluation Quality Assurance (EQA) as a tool to guide the field toward systemic solutions rather than inadvertently propagating issues in evaluation. Overall, this guide aims to provide a roadmap for higher quality data and better grounded evaluations, ultimately improving research in misinformation detection. All datasets and other artifacts are available at misinfo-datasets.complexdatalab.com.
@@ 1. Introduction
Finally, our findings come together in our recommendation for a new practice of Evaluation Quality Assurance (EQA), which offers a tool to improve data selection and evaluation methodologies. In summary, we present a guide to misinformation datasets, including:
@@
  • Recommended practices of EQA, for research proposing misinformation detection methods. Informed by our findings of severe quality issues in both data and evaluation procedures, such as how simple metrics like accuracy and F1 may be obsolete in this domain, we propose that assurance of evaluation quality is an indispensable component for future research here.
@@ 3. Data Quality
Shortcut learning is a serious barrier to predictive systems working in the real world (Geirhos et al., 2020). In this section, we assess datasets’ potential for teaching algorithms spurious keyword and temporal correlations. These are key considerations when training and evaluating supervised methods. We then assess whether the inputs are making sufficiently complete and unambiguous claims for it to be feasible to predict their veracity at all, which is critical to the evaluation of all types of methods. We provide the automated tools that enable this analysis through our website as the Complex Data Lab Dataset Quality Assessment (CDL-DQA).
@@ 3.3. Feasibility
We first aggregate over datasets, presenting the individual results of our 9 human and AI annotators in Figure 4. Our results show universal agreement that without an evidence retrieval (search) system, at least half of claims in these 29 datasets cannot be validly assessed for veracity. This suggests that such methods are often being evaluated on impossible tasks, and there is a severe risk that evaluation comparing such methods is determining not the best generalizable predictor but the best shortcut learner. We recommend that any proposed methods in this category conduct stringent Evaluation Quality Assurance (see Section 5), and suggest that applications requiring even slightly out of distribution predictions should ground them in evidence.
@@
This has 3 main implications. First, there is an urgent need for higher quality claims datasets. Second, supplementary information beyond claim text, such as claim dates, authors, and other additional context may help alleviate this problem. However, it is critical to assess how the specific information a predictive method processes impacts the feasibility of the data it is being trained and evaluated on, to confirm that performance margins between methods represent real progress and not progress in predicting noise. Thus, third, we propose that even in the more favorable retrieval-augmented setting, Evaluation Quality Assurance (Section 5) is paramount.
@@ 4.2. The Flaw in Current Metrics
Label $\neq$ Prediction Label $=$ Prediction
AI rationale is: Valid Invalid
@@
5. Evaluation Quality Assurance
@@
We have highlighted multiple ways datasets and evaluation procedures could deliver results that will fail to translate to the real world. While the information here directly reveals some pitfalls, we suspect that future research—with new datasets, new methods, and new needs—will encounter similar problems. To improve the validity of conclusions in our field in a systematic and future-proof way, we propose Evaluation Quality Assurance (EQA) as a key component of an improved research methodology.
@@
We draw inspiration from the common practices for systematically and transparently documenting dataset publications, such as “datasheets for datasets” (Gebru et al., 2021), and suggest that similar attention should be given to the usage of datasets. However, it is important to ensure data and evaluation quality not just as a lengthy post-hoc appendix, but as an integral part of the research process and a key foundation of every paper’s core arguments. Therefore, we propose including EQA as a scientifically essential step within the methodology. This means augmenting the standard descriptive discussion of data used with a critical analysis of it, explaining what steps and quality assessments were conducted to assure that the data is suitable to prove the experimental conclusions. We suggest that dedicating even just one paragraph to this, rather than taking quality of “standard” datasets for granted, could dramatically improve the validity of our empirical results and the rate of real progress in the field. Conversely, current practices risk misleading evaluations, generalization failures, and immeasurable time and effort dedicated to measurements of noise.
We make two key recommendations for Evaluation Quality Assurance.
@@
  • Quality Assessment: analyze the data and evaluation processes themselves, not just their results, to assure they are suitable to test performance and generalization of a method. One basic analysis, that can be nearly universally recommended, is to qualitatively inspect at least 25 random examples—input, output, and ground truth—where the method made correct predictions, and 25 where it made incorrect ones. This data can be shared for reproducibility, and provides a basic sanity check. This simple step could already have a high impact; for example, the feasibility and categorical metrics issues we have reported would be detected by such a sanity check in the majority of the datasets that we examined. Furthermore, we also recommend papers perform quantitive assessments of their data and evaluation, such as (but not limited to) the keyword, temporal, and feasibility assessments discussed in prior sections and available in our CDL-DQA toolkit.
  • Assurance Limitations: report limitations in evaluation quality assurance (EQA), not only in terms of dataset size or experimental scope but also in terms of quality issues that the EQA could not rule out. Key to this is considering how a method would be applied in the real world—and how the evaluation might not be fully capturing the real-world conditions. If, for instance, a method is intended to handle diverse contexts or domains of misinformation, but was only tested on data from a single context (e.g., one country) or domain (e.g., politics), this should be stated as a limitation so that future research can address it. Explicitly acknowledging such gaps strengthens scientific rigor and ensures that claims are appropriately supported by the evidence provided.
@@
In this way, we treat assuring evaluation quality as a crucial step for validating the research process. For authors, this will help reinforce the demonstrated rigour of the experiments and the validity of proposed methods. It will also prevent the rejection of strong methods due to unrealistically high baseline settings caused by spurious correlations or other evaluation issues. For reviewers, this will produce more grounding (or expose the lack thereof) for the experiments, as well as a framework for recommendations when evaluation is insufficiently justified and needs better validation. Thus, we propose that adopting EQA can efficiently and substantively mitigate propagating and compounding errors, that arise in this domain from current practices of copy-pasting evaluation from past literature.